# A formal implementation of Behavior Trees to act in robotics

### v1

@@ Abstract
Behavior Trees (BT) are becoming quite popular as an Acting component of autonomous robotic systems. We propose to define a formal semantics to BT by translating them to a formal language which enables us to perform verification of programs written with BT, as well as runtime verification while these BT execute. This allows us to formally verify BT correctness without requiring BT programmers to master formal language and without compromising BT most valuable features: modularity, flexibility and reusability. We present the formal framework we use: Fiacre, its langage and the produced TTS model; Tina, its model checking tools and Hippo, its runtime verification engine. We then show how the translation from BT to Fiacre is automatically done, the type of formal LTL and CTL properties we can check offline and how to execute the formal model online in place of a regular BT engine. We illustrate our approach on two robotics applications, and show how BT could benefit of other features available in the Fiacre formal framework (state variables, time, etc).
@@ 1 Introduction and motivation
Considering than more and more robotics applications, using BT, may be deployed in critical application (autonomous drones or vehicles, etc) or in environment with human (service robots), we need to be able to prove some safety properties on the BT, and to trust their execution will remain faithful to the programmer intentions. For this we believe we need to harness some formal models to the BT language and its execution engine to enable some formal offline and online validation and verification of BT.
@@ 2.1 BT in robotics
BT are praised for their modularity, readability, scalability, flexibility and robustness, and supposedly being easy to debug and test. Even if these are questionable and somewhat subjective, one cannot deny their rising success and interest in robotics for autonomous navigation, human-robot interaction, manipulation tasks and multi-agent systems.
@@ 2.2 BT and formal models
The authors of Colledanchise et al. (2021) focus on the formalisation of the execution context of BT to be able to perform runtime verification. They propose channel systems to model the “surroundings” of the BT and then to check at runtime that some specifications, written in the SCOPE langage, are satisfied while the BT executes. So the approach, which is not limited to BT, provides a very strong and formal execution framework sitting between the BT and the robot, behave like a safety bag. The battery example they present has some similarity with the one we deploy on our UAV in section 6.1.
@@ 3 A Formal Framework for Offline and Runtime Verification: The Fiacre Language, Models, and Tools
While this paper does not aim to exhaustively present the formal framework we use, some terminology and explanations are essential for clarity and make the paper self contained. Readers interested in more details may consult the specific papers and websites cited below.  [^2] Note that a similar Fiacre presentation can be found in this paper Ingrand (2024a) (from the same author). We include it almost as is in this paper as to make the paper self contained, nevertheless, If the reviewers believe this section should be shorten and replace by a pointer to the other paper, this is perfectly fine for us.
@@ 3.1 Terminology, Models, Languages, and Tools
Berthomieu and Diaz (1991) are an extension of traditional Petri nets where each transition has an associated time interval (typically $[0,\infty)$) specifying the time range within which a sensitized transition can be fired.
@@
short for ”TIme Petri Net Analyzer,” is a toolkit for editing, simulating, and analyzing Petri nets, Time Petri nets, and TTS. Within this toolkit, sift and selt enable the construction of reachable state sets and the verification of LTL properties. [^3] https://projects.laas.fr/tina/index.php
@@
extends the Fiacre language, making models “executable” by adding Event Ports and Tasks linked to C/C++ functions.
@@ 3.2 Fiacre Semantics
Following these specifications, a component is defined by placing three process instances in parallel (line 1.56) and linking them through two ports (line 1.54). This example is simple by design, though the Fiacre language supports complex data types, bidirectional ports, local and global variables, conditions, switch/case statements, transition guards, and function calls (internal to Fiacre or external in C/C++ code) for advanced computation. More complex Fiacre specifications can be found in A and B.
@@
67
68property cannot_receveice_tc is absent comp_tc/3/state received_tc
@@ 3.4 H-Fiacre Runtime Extensions
are also defined in the preamble (see line 2.2), associating a task (in this case, report_triplec) with a C/C++ function (here c_report_triple_click), which is called asynchronously upon a start (see line 2.17). This enables the corresponding sync (line 2.21) once the C/C++ function completes. Values can be passed to the task at call time and returned when it completes.
@@
2task report_triplec () : nat is c_report_triple_click // The C/C++ function called by this task
3
4process detect_triple_click [triple_click:sync] is
5// this process is exactly the same than in the regular Fiacre version
6// only the click port is now an event port
7
8process triple_click_receiver[triple_click:sync] is
9states waiting_tc, received_tc, sync_report
10var ignore : nat
11
12from waiting_tc
@@
15
16from received_tc // show an example of an external call
@@
19
20from sync_report
@@
23
24component comp_tc is
25port triple_click:sync
26
27par * in
@@
29|| triple_click_receiver[triple_click]
30end
31
32comp_tc
@@ 3.5 Runtime (Online) Verification
Moreover, if runtime behavior diverges from expectations, you can debug it as you would a typical program. From a formal perspective, the traces (or behavior patterns) of the H-Fiacre version (also called the Tina version) are contained within those of the Fiacre model (also called the Hippo version), ensuring consistency between the offline model and its runtime counterpart.
@@ 4 The mapping of BT in Fiacre
Listing 4: A simple drone survey BT. Note that for historical reasons and as we are reusing some of our existing tools, we do not use an XML syntax but rather a Lisp like syntax (called the .btf format), but this has no consequence on the content and the interpretation of BT (See Appendix C, Listing 10p10 and 11p11 for an example in both format).
@@ 4.2 Condition Node
The Condition node is the simplest node, it only returns success or failure. The Tina version will just return either values (See Figure 6p6), while the Hippo version will call the external C/C++ function which performs the test and return its value (See Figure 7p7). The C/C++ function is expected to be fast and should not delay execution unnecessarily. Note that in our formalism, one can pass arguments to the C/C++ call. This is very much welcome to make the BT language more expressive and to some extends to have these values available in the formal model.
@@ 4.3 Action Node
The Action node is slightly more complex, as it can also return running. Indeed, actions may take some time and not return success or failure right away. To preserve the reactivity of the overall execution, and keep the tick short, it is often advised to return running while the action is still executing in its own thread. As a consequence, it means that an Action node may be halted (i.e. at some point, it is in a running state but its parent node wants to halt it). In the Tina version (See Figure 8p8), all the return values are possible, while in the Hippo version (See Figure 9p9), a Fiacre task is started calling the C/C++ action_task, and subsequent ticks will return running until the C/C++ action_task finishes and returns success or failure. An external C/C++ action_halt is also defined and is called if the action must be halted (it then returns failure). While action_task may take some time in its own thread, the action_halt is expected to be fast. In an Action node too, one may define and pass some arguments to the C/C++ function. Listing 8p8 in Appendix A show the Fiacre code for the takeoff action.
@@ 4.4 Sequence Nodes
  • success, if it is the last child to success, we return success, otherwise, we proceed to the next child $Child_{\texttt{next\_seq+1}}$.
@@
There exist variants of Sequence: ReactiveSequence, SequenceWithMemory. They alter the way the sequence is ticked after failure or running are returned. For example ReactiveSequence always restarts the sequence from the beginning after one child has returned running. The goal here is not to list all the variants, just to give an hint on how these are transformed in Fiacre. Of course, the generated Fiacre code implements the proper semantics for each variant by properly setting the next_seq value.
Note that the Tina and the Hippo versions of the BT Fiacre process are exactly the same.
@@ 4.5 Fallback Nodes
Figure 11: The Fiacre process modeling the Fallback node.
@@ 4.6 Parallel Nodes
The Parallel node specifies how many children m out of all n children must succeed for the Parallel node to succeed (with ParallelAll they must all succeed: n = m). From this, we see that the implementation just ticks all the node, and then keep track of how many return failure, success or running (See Figure 12p12). If any final condition for success or failure is met (enough BTnode[child1].rstatus = success)888(enough… success) is true when success >= m or (enough BTnode[childn].rstatus = failure)999(enough… failure) is true when failure > n - m, then it proceeds to the corresponding state, otherwise, this node returns running. Note that Parallel nodes can lead to children being halted (if the condition for success or failure are met while some children are still running).
Figure 12: The Fiacre process modeling the Parallel node.
@@ 4.8.2 Ticks and BT root
The BT root is also responsible for keeping the BT alive. What should it do when its child return running, success or failure? Some implementation keep running on success, other do not. To clarify the model we produce, the root BT keeps running while its child returns running, and stops when it returns success or failure (e.g., in the drone experiment we present in Section 6.1, this is the expected behavior).
@@ 5 Deploying, Model Checking and Running the Fiacre BT
As shown on Figure 3p3 page 3 both Fiacre models (offline and online) are compiled in their respective TTS (i.e., a Time Petri Net with data) with the frac compiler.
@@ 5.1 Offline formal verification of BT
One can also check more complex properties. For example that if we reach a particular state, then we will eventually reach another one.
We also modelled some examples found in some papers presented in the section 2.1.
In Biggar and Zamani (2020) the authors present an example of a Mars rover using unfolded solar panel to charge its battery, but should fold them when there is a storm. The same example is studied in  Wang et al. (2024) with the BT transformed to BIP on which they check the same properties. The equivalent BT in the .btf formalism can be found in Appendix D, Listing 12p12. After translation to Fiacre, we can build the set of reachable states:
@@
Similarly, in Wang et al. (2024) they develop an example with a train having a non nul speed while a door is open, we modelled it and reach the same formal proof results.
@@ 5.2 Online (or runtime) verification
As seen on Figures 4p4 and 14p14, our BT show changing color while executing: white, the node has not yet been visited; dark blue, the tick is currently in this node; light blue, the tick has been passed to child(ren) node(s) below and the node is waiting for the children to return; yellow, the node has returned running; green, the node has returned success; red, the node has returned failure; purple, the node was running and has been halted by one of its parent nodes; pink, the node has been instructed to halt itself (and its running branches).
@@ 6.1 An UAV surveying an area
We program an UAV to perform a survey mission with a BT. The functional layer of this experiment (Figure 14p14) has already been presented in Dal Zilio et al. (2023), but suffice to say that it provides robust localization, navigation, flight control and allows us to command the drone and its camera. It is deployed using the G en oM specification language (which also maps in the same formal framework to validate and verify the functional components Dal Zilio et al. (2023)).
@@ 6.1.1 Offline verification results
which takes around 1 hour and 16 minutes to build on an Intel(R) Xeon(R) Silver 4110 CPU @ 2.10 GHz with 256 GB of memory. The resulting .ktz has $49196302$ classes, $49145735$ markings, $152$ domains, $267702880$ transitions.
@@
Among the default proven properties, somes are still interesting to comment. For example, despite the ParallelAll on the start_drone and the start_camera, they cannot be halted (because the ParallelAll has :halt 0). Unlike the BT children of the ReactiveSequence which has :halt 1, they can all be halted (except the first branch) [^13] As only a branch returning running on the previous tick may be halted if one branch on its left fails, hence, never the first branch..
@@
ltl ($\Box$ (drone/5/value (battery=Critical)) $\Rightarrow$
@@
ltl ($\Box$ (drone/12/value (battery=Low)) $\Rightarrow$
@@
ltl ($\Box$ (drone/16/value (btnode[localization_ok_btn18].rstatus = failure)) $\Rightarrow$
@@
(drone/5/value (battery=Critical)) leadsto (drone/8/state tick_node) within [0,2]
@@ 6.2.1 Runtime
All the Nav2 BT have been translated to .btf and run with Hippo and simulated ROS2 actions. One of the features we added and we believe goes beyond the “nice graphical touch” aspect, is to dynamically color the BT nodes while they execute. Of course, there are text traces of the Hippo engine running the BT, but being able to follow the execution ticks as well as the last returned status for each BT is rather informative. Some videos of these runs and the BT animation are available here: https://redmine.laas.fr/projects/bt2fiacre/pages/index#playing-with-ros2-nav2-bt. Figure 17p17 shows the Nav2 BT (See Listing 11p11 and Figure 18p18) executed by Hippo.
Figure 17: The screen dump of the Nav2 BT (See Listing 11p11 and Figure 18p18) executing.
@@ 7.1 Limits
An area where our approach would suffer, is when BT are dynamically modified, or transformed. For example in Section 2.1 we consider BT used for planning. In these approaches, BT are being modified on the fly, clearly, we could not perform model checking on the fly, yet assuming the dynamic of the planner remains slow (below a second) we could still synthesize the model, compile it and jump in its execution on the fly.
Although all our examples are based on one BT, absolutely nothing prevent us from having multiple BT executing together. yet from a verification point of view, this would probably lead to intractable model as models running in parallel tend to multiply the size of their reachable states set, to account for all the possible transition interleavings.
This brings us to the main limiting factor of our approach. Offline verification with model checking may lead to large intractable reachable states set. This is a well known limit of these approaches and we do not have any magic bullet. Yet we have seen that we can verify reasonably complex robotics BT skills, that we can prove “individual” BT node behavior (e.g. the ones added in Nav2), and that we can also handle complex properties on missions written with BT.

### v3

@@ Abstract
Behavior Trees (BT) are becoming quite popular as an Acting component of autonomous robotic systems. We propose to define a formal semantics to BT by translating them to a formal language which enables us to perform verification of programs written with BT, as well as runtime verification while these BT execute. This allows us to formally verify BT correctness without requiring BT programmers to master formal languages and without compromising BT most valuable features: modularity, flexibility and reusability. We present the formal framework we use: Fiacre, its language and the produced TTS model; Tina, its model checking tools and Hippo, its runtime verification engine. We then show how the translation from BT to Fiacre is automatically done, the type of formal LTL and CTL properties we can check offline and how to execute the formal model online in place of a regular BT engine. We illustrate our approach on two robotics applications, and show how BT can be extended with state variables, eval nodes, node evaluation results and benefit of other features available in the Fiacre formal framework (e.g., time).
@@ 1 Introduction and motivation
Considering that more and more robotics applications, using BT, may be deployed in critical applications (autonomous drones or vehicles, etc) or in an environment with human (service robots), we need to be able to prove some safety properties on the BT, and to trust their execution will remain faithful to the programmer’s intentions. For this we believe we need to harness some formal models to the BT language and its execution engine to enable some formal offline and online validation and verification of BT.
@@ 2.1 BT in robotics
BT are praised for their modularity, readability, scalability, flexibility, robustness, and supposedly being easy to debug and test. Even if these are questionable and somewhat subjective, one cannot deny their rising success and interest in robotics for autonomous navigation, human-robot interaction, manipulation tasks and multi-agent systems.
@@ 2.2 BT and formal models
The authors of Colledanchise et al. (2021) focus on the formalisation of the execution context of BT to be able to perform runtime verification. They propose channel systems to model the “surroundings” of the BT and then to check at runtime that some specifications, written in the SCOPE language, are satisfied while the BT executes. So the approach, which is not limited to BT, provides a very strong and formal execution framework sitting between the BT and the robot, behaving like a safety bag. The battery example they present has some similarity with the one we deploy on our UAV in section 6.1.
@@ 3 A Formal Framework for Offline and Runtime Verification: The Fiacre Language, Models, and Tools
While this paper does not aim to exhaustively present the formal framework we use, some terminology and explanations are essential for clarity and make the paper self contained. Readers interested in more details may consult the specific papers and websites cited below.  [^2] Note that a similar Fiacre presentation can be found in this paper Ingrand (2024a) (from the same author). We include it almost as is in this paper as to make the paper self contained, nevertheless, If the reviewers believe this section should be shortened and replaced by a pointer to the other paper, this is perfectly fine for us.
@@ 3.1 Terminology, Models, Languages, and Tools
Berthomieu and Diaz (1991) are an extension of traditional Petri nets where each transition has an associated time interval (typically $[0,\infty)$) specifying the time range within which an enabled transition can be fired.
@@
(short for ”TIme Petri Net Analyzer”) is a toolkit for editing, simulating, and analyzing Petri nets, Time Petri nets, and TTS. Within this toolkit, sift and selt enable the construction of reachable state sets and the verification of LTL properties. [^3] https://projects.laas.fr/tina/index.php
@@
adds Event Ports and Tasks linked to C/C++ functions, to make the Fiacre models “executable”.
@@ 3.2 Fiacre Semantics
Following these specifications, a component is defined by placing three process instances in parallel (line 1.56) [^8] Listing lines are referenced with the $<$listing number$>$.$<$line number$>$, example: 1.56 is Listing 1, line: 56. and linking them through two ports (line 1.54). This example is simple by design, though the Fiacre language supports complex data types, bidirectional ports, local and global variables, conditions, switch/case statements, transition guards, and function calls (internal to Fiacre or external in C/C++ code) for advanced computation. More complex Fiacre specifications can be found in A and B.
@@
67// in the next property comp_tc/3/state designates the state in the 3rd process of
68property cannot_receveice_tc is absent comp_tc/3/state received_tc // the comp_tc component
@@ 3.4 H-Fiacre Runtime Extensions
are also defined in the preamble (see line 2.3), associating a task (in this case, report_triplec) with a C/C++ function (here c_report_triple_click), which is called asynchronously upon a start (see line 2.18). This enables the corresponding sync (line 2.22) once the C/C++ function completes. Values can be passed to the task at call time and returned when it completes.
@@
2task report_triplec () : nat is // declare the task and
3 c_report_triple_click // the C/C++ function called by this task
4
5process detect_triple_click [triple_click:sync] is
6// this process is exactly the same than in the regular Fiacre version
7// only the click port is now an event port
8
9process triple_click_receiver[triple_click:sync] is
10states waiting_tc, received_tc, sync_report
11var ignore : nat
12
13from waiting_tc
@@
16
17from received_tc // show an example of an external call
@@
20
21from sync_report
@@
24
25component comp_tc is
26port triple_click:sync
27
28par * in
@@
30|| triple_click_receiver[triple_click]
31end
32
33comp_tc
@@ 3.5 Runtime (Online) Verification
Moreover, if runtime behavior diverges from expectations, you can debug it as you would with any programs. From a formal perspective, the possible traces of the H-Fiacre version (also called the Hippo version) are contained within those of the Fiacre model (also called the Tina version), ensuring consistency between the runtime model and its offline counterpart.
@@ 4 The mapping of BT in Fiacre
Listing 4: A simple drone survey BT. Note that for historical reasons and as we are reusing some of our existing tools, we do not use an XML syntax but rather a Lisp like syntax (called the .btf format), but this has no consequence on the content and the interpretation of BT (See Appendix E, Listing 10p10 and 11p11 for an example in both formats).
@@ 4.2 Condition Node
The Condition node is the simplest node, it only returns success or failure. The Tina version will just return either values (See Figure 5p5), while the Hippo version will call the external C/C++ function which performs the test and return its value (See Figure 6p6). The C/C++ function is expected to be fast and should not delay execution unnecessarily. Note that in our formalism, one can pass arguments to the C/C++ call. This is very much welcome to make the BT language more expressive and to some extent to have these values available in the formal model.
@@ 4.3 Action Node
The Action node is slightly more complex, as it can also return running. Indeed, actions may take some time and not return success or failure right away. To preserve the reactivity of the overall execution, and keep the tick short, it is often advised to return running while the action is still executing in its own thread. As a consequence, it means that an Action node may be halted (i.e. at some point, it is in a running state but its parent node wants to halt it). In the Tina version (See Figure 7p7), all the return values are possible, while in the Hippo version (See Figure 8p8), a Fiacre task is started calling the C/C++ action_task, and subsequent ticks will return running until the C/C++ action_task finishes and returns success or failure. An external C/C++ action_halt is also defined and is called if the action must be halted (it then returns failure). While the action_task may take some time in its own thread, the action_halt is expected to be fast. In an Action node too, one may define and pass some arguments to the C/C++ function. Listing 8p8 in Appendix A shows the Fiacre code for the takeoff action.
@@
Note that only the Action and Condition nodes have a slightly different Tina (offline) and Hippo (online) versions. For all the other BT nodes, the model is strictly the same.
@@ 4.4 Sequence Nodes
  • success, if it is the last child to succeed, we return success, otherwise, we proceed to the next child $Child_{\texttt{next\_seq+1}}$.
@@
There exist variants of Sequence: ReactiveSequence, SequenceWithMemory. They alter the way the sequence is ticked after failure or running are returned. For example ReactiveSequence always restarts the sequence from the beginning after one child has returned running. The goal here is not to list all the variants, just to give a hint on how these are transformed in Fiacre. Of course, the generated Fiacre code implements the proper semantics for each variant by properly setting the next_seq value.
@@ 4.6 Parallel Nodes
The Parallel node specifies how many children m out of all n children must succeed for the Parallel node to succeed (with ParallelAll they must all succeed: n = m). From this, we see that the implementation just ticks all the node, and then keep track of how many return failure, success or running (See Figure 15p15). If any final condition for success or failure is met (enough BTnode[child1].rstatus = success)999(enough… success) is true when success >= m or (enough BTnode[childn].rstatus = failure)101010(enough… failure) is true when failure > n - m, then it proceeds to the corresponding state, otherwise, this node returns running. Note that Parallel nodes can lead to children being halted (if the condition for success or failure are met while some children are still running).
@@ 4.8.2 Ticks and BT root
The BT root is also responsible for keeping the BT alive. What should it do when its child returns running, success or failure? Some implementations keep running on success, others do not. To clarify the model we produce, the root BT keeps running while its child returns running, and stops when it returns success or failure (e.g., in the drone experiment we present in Section 6.1, this is the expected behavior).
@@ 5 Deploying, Model Checking and Running the Fiacre BT
As shown on Figure 3p3 both Fiacre models (offline and online) are compiled in their respective TTS (i.e., a Time Petri Net with data) with the frac compiler.
@@ 5.1 Offline formal verification of BT
One can also check more complex properties. For example, if we reach a particular state, then we will eventually reach another one.
We also modelled some examples found in some papers presented in Section 2.1.
In Biggar and Zamani (2020) the authors present an example of a Mars rover using unfolded solar panels to charge its battery, but should fold them when there is a storm. The same example is studied in  Wang et al. (2024) with the BT transformed to BIP on which they check the same properties. The equivalent BT in the .btf formalism can be found in Appendix F, Listing 12p12. After translation to Fiacre, we can build the set of reachable states:
@@
Similarly, in Wang et al. (2024) they develop an example with a train having a non null speed while a door is open, we modelled it and reach the same formal proof results.
@@ 5.2 Online (or runtime) verification
As seen on Figures 16p16 and 11p11, our BT show changing color while executing: white, the node has not yet been visited; dark blue, the tick is currently in this node; light blue, the tick has been passed to child(ren) node(s) below and the node is waiting for the children to return; yellow, the node has returned running; green, the node has returned success; red, the node has returned failure; purple, the node was running and has been halted by one of its parent nodes; pink, the node has been instructed to halt itself (and its running branches).
@@ 6.1 An UAV surveying an area
We program an UAV to perform a survey mission with a BT. The functional layer of this experiment (Figure 11p11) has already been presented in Dal Zilio et al. (2023), but suffices to say that it provides robust localization, navigation, flight control and allows us to command the drone and its camera. It is deployed using the G en oM specification language (which also maps in the same formal framework to validate and verify the functional components Dal Zilio et al. (2023)).
@@ 6.1.1 Offline verification results
which takes around 1 hour and 16 minutes to build on an Intel(R) Xeon(R) Silver 4110 CPU @ 2.10 GHz with 256 GB of memory. The resulting .ktz has $49196302$ classes, $49145735$ markings, $152$ domains, $267702880$ transitions. In this context, a marking is a particular set of states and values for all the processes and variables in the system. A class is a state extended with timing information on the enabled transitions (therefore we can have several classes with the same marking).
@@
Among the default proven properties, some are still interesting to comment on. For example, despite the ParallelAll on the start_drone and the start_camera, they cannot be halted (because the ParallelAll has :halt 0). Unlike the BT children of the ReactiveSequence which has :halt 1, they can all be halted (except the first branch) [^14] As only a branch returning running on the previous tick may be halted if one branch on its left fails, hence, never the first branch..
@@
ltl ($\Box$ ((drone/5/value (battery=Critical)) and (drone/5/state done)) $\Rightarrow$
@@
ltl ($\Box$ ((drone/12/value (battery=Low)) and (drone/5/state done)) $\Rightarrow$
@@
ltl ($\Box$ ((drone/16/value (btnode[localization_ok_btn18].rstatus = failure) and (drone/16/state done))) $\Rightarrow$
@@
((drone/5/value (battery=Critical))and (drone/5/state done)) leadsto (drone/8/state tick_node) within [0,2]
@@ 6.2.1 Runtime
All the Nav2 BT have been translated to .btf and run with Hippo and simulated ROS2 actions. One of the features we added and we believe goes beyond the “nice graphical touch” aspect, is to dynamically color the BT nodes while they execute. Of course, there are text traces of the Hippo engine running the BT, but being able to follow the execution ticks as well as the last returned status for each BT is rather informative. Some videos of these runs and the BT animation are available here: https://redmine.laas.fr/projects/bt2fiacre/pages/index#playing-with-ros2-nav2-bt. Figure 17p17 shows the Nav2 BT (See Listing 11p11) executed by Hippo.
@@ 7.1 Limits
An area where our approach would suffer, is when BT are dynamically modified, or transformed. For example in Section 2.1 we consider BT used for planning. In these approaches, BT are being modified on the fly, clearly, we could not perform model checking on the fly, yet assuming the dynamic of the planner remains slow (below one second) we could still synthesize and compile the model (both are almost instantaneous), and jump in its execution on the fly.
Although all our examples are based on one BT, absolutely nothing prevents us from having multiple BT executing together. Yet from a verification point of view, this would probably lead to an intractable model as models running in parallel tend to multiply the size of their reachable states set, to account for all the possible transition interleavings.
This brings us to the main limiting factor of our approach. Offline verification with model checking may lead to large intractable reachable states set. This is a well known limit of these approaches and we do not have any magic bullet. It is hard to predict what will be the size of the reachable states set for a given BT. Some very large BT may still produce a rather small reachable states set, while a simple BT (e.g., with a lot of parallelism, loops, etc) are untractable. Yet we have seen in our examples that we can verify reasonably complex robotics BT skills, that we can prove “individual” BT node behavior (e.g. the new control nodes added in Nav2), and that we can also handle complex properties on missions written with BT. For the models too complex to be verified offline, we can still use the complete online version with some added monitors, or work on an abstracted model for offline verification (but we would then loose the offline/online equivalence).